Integration Procedures (Integration by Parts)


∫csc^(-2)x dx


1
Expert's answer
2021-04-29T16:56:23-0400

"I=\int{csc^{-2}x}dx=\int{sin^2x dx}=-\cos{x}\sin{x}-"


"-\int{(-\cos{x})\cos{x}dx}="


"=-\cos{x}\sin{x}+\int{\cos^2{x}dx}="


"=-\cos{x}\sin{x}+\int{(1-\sin^2{x})dx}="


"=-\cos{x}\sin{x}+x-\int{\sin^2{x}dx}="


"=-\cos{x}\sin{x}+x-I"


"2I=-\cos{x}\sin{x}+x"


"I=\frac{1}{2}(x-\cos{x}\sin{x})"